OpenStudy (anonymous):

Write an equation of the line that passes through the point (1,2) and has the slope, m = 2.

OpenStudy (anonymous):

i know its y-y1=m(x-x1)

OpenStudy (anonymous):

@Princer_Jones

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

@jojo4eva

OpenStudy (anonymous):

@Here_to_Help15

OpenStudy (jojo4eva):

y-2=2(x-1) y-2=2x-2 Add 2 to both sides to isolate variable y and you answer is: y=2x

OpenStudy (jojo4eva):

So the answer is y=2x

OpenStudy (anonymous):

really

OpenStudy (jojo4eva):

yup what did u get